OpenStudy (anonymous):

A toy rocket is launched upward from ground level with an initial velocity of 42 meters per second. Its height is a function of time, given by h=-4.9^2+42t, where h represents the height in meters and t represents the time in seconds. a. The maximum height reached by the rocket is? b. The minimum height reached by the rocket is?

OpenStudy (anonymous):

the second coordinate of the vertex the first coordinate is \(-\frac{b}{2a}\) which in your case is \[-\frac{42}{2\times -4.9}\]

OpenStudy (anonymous):

i get \[\frac{30}{7}\] as the first coordinate not so pretty but you can plug it in to get the second coordinate ,which is your answer

OpenStudy (anonymous):

the minimum is zero, when it hits the ground doh

OpenStudy (anonymous):

how do I figure out the maximum?

OpenStudy (danjs):

oh same asker too. i see what you did there